OpenStudy (anonymous):

Initial investment:$750 Annual rate:? Time to double: 7 and 3/4 yrs Amount after 10 yrs:? Using the formula A=Pe^rt, find the annual rate, and amount after 10 years.. thanks so much!!! please explain!!

OpenStudy (slaaibak):

A = Accumulated amount P = Start amount r = annual rate t = time. This should help you solve the question.

OpenStudy (slaaibak):

You'll learn a lot more by doing it :)

OpenStudy (anonymous):

A=750e^r(10) thats how far i got.. dont know what do to now

OpenStudy (slaaibak):

You first need to solve r. They tell you it doubles in 7 and 3/4 years. They give you P (the initial value) and they also give you A (indirectly) So you have A, t, and P. solve r. Once you have r, you'll be able to solve the second part of the question

OpenStudy (anonymous):

so i just dont know what A is then...?

OpenStudy (slaaibak):

To get the A after 10 years, you do that which you said previously: A=750e^r(10) BUT You don't know what r is yet. So the only way to solve r, is to find out at which interest rate the 750 doubled in 7 and 3/4 years. To find this, plug those value into the formula. After you find the r, you can find the amount after 10 years

OpenStudy (anonymous):

oh so 1500=750e^r(7and3/4)?

OpenStudy (slaaibak):

yes!

OpenStudy (anonymous):

so the answer to this is the rate?

OpenStudy (slaaibak):

Yep. r is the rate

OpenStudy (anonymous):

so after this would the A be 1500 as well???

OpenStudy (slaaibak):

No. r is a constant. So, once you get r, you plug it into this: A=750e^r(10)

OpenStudy (anonymous):

ok so i got 9.22 A=750e^9.22(10) now do you just divide 750, then do the LN?

OpenStudy (anonymous):

or you just times everything..i think

OpenStudy (slaaibak):

How did you get 9.22? You sure about that? Just recheck, show me what you did, something went wrong

OpenStudy (anonymous):

1500=750e^r(7and3/4) divide 750 2=e^r(7 and 3/4) take Ln LN 2=r(7and 3/4) divide by 7.75 Ln 2/7.75 =0.0894383

OpenStudy (slaaibak):

Yes, thats correct. Now remember, when you use that in the formula, you should use 0.0894383. It is 8.94383%, but if used in calculation, it should be 0.0894383. Now you can solve for A!

OpenStudy (anonymous):

wait so do i use 8.94383? or the other?

OpenStudy (anonymous):

so it would be A=750e^0.0894383(10)??

OpenStudy (slaaibak):

Yep

OpenStudy (anonymous):

Thank you so much!! You are AMAZING!! This helped tons!!

OpenStudy (slaaibak):

Shot, glad to help! Shout out if you need anything else.

OpenStudy (anonymous):

Will do thanks:)